OpenStudy (anonymous):

I need help with variables on both sides. If you could show me how to do it without telling me the answer that would be great. 7r - 7 = 2r + 18

OpenStudy (anonymous):

Start by adding together like terms. Pick either the numbers or the variables to add/subtract from one side. Do the same to both sides. Ex. 80 + 9x= 90+ 7x I want to subtract eighty--->9x=10+7x

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

Then....you have 9x=10+7x. In this case, your like terms are the variables. So you should subtract 7x from both sides to get 2x= 10. 9x-7x=10+7x-7x 2x=10 Now to solve for x, you divide out the two. 2x/2=10/2 x=5 Can you do this with your equation?

OpenStudy (anonymous):

Ok. Thanks that helps.